# Exactly solvable dissipative dynamics and one-form strong-to-weak spontaneous symmetry breaking in interacting two-dimensional spin systems

### v1

@@ II.5 Strong and weak symmetries
Unitary symmetries of open quantum systems come in two flavors: strong and weak [123] (sometimes also called exact and average, respectively), the properties of which we now briefly review. Let $S$ be a unitary operator in the physical Hilbert space. $S$ is a weak symmetry of $\mathcal{L}$ if $\mathcal{L}(S\rho S^{\dagger})=S\mathcal{L}(\rho)S^{\dagger}$, for any operator $\rho$. Introducing the superoperator $\mathcal{S}$ that acts by conjugation with $S$ (i.e., the adjoint action of $S$ on the space of operators), $\mathcal{S}(\rho)=S\rho S^{\dagger}$, a weak symmetry is equivalent to the superoperator commutation relation $[\mathcal{L},\mathcal{S}]=0$. On the other hand, $S$ is a strong symmetry if $S$ commutes with all elements of the set $\{H,L_{j},L_{j}^{\dagger}\}$ ($j\in\{1,\dots,N\}$). It follows that the Lindbladian superoperator commutes with both $\mathcal{S}_{+}$ and $\mathcal{S}_{-}$, where $\mathcal{S}_{+}$ is the left action of $S$, $\mathcal{S}_{+}(\rho)=S\rho$, and $\mathcal{S}_{-}$ the right action, $\mathcal{S}_{-}(\rho)=\rho S^{\dagger}$. Clearly, a strong symmetry is also weak, but the converse is not true. A simple example of a weak symmetry that is not strong (and that occurs in our model) is if $S$ commutes with $H$ but anticommutes with (some of) the jump operators.
@@ II.6 Symmetries of the gamma-matrix Lindbladian
Not all $W_{\mathcal{C}}$, $\widetilde{W}_{\mathcal{C}}$ and $\mathcal{W}_{\mathcal{P}}^{\mathrm{ch}}$ are independent. As worked out in detail in App. B, a minimal set—which labels the symmetry sectors—is $\{W_{p},W_{\ell},\mathcal{W}_{q}\}_{p,\ell,q}$. Here, $W_{p}$ are $W_{\mathcal{C}}$ for $\mathcal{C}=p$, with $p$ the elementary plaquettes (or faces) of the graph. If the graph is embedded on a closed surface, the total flux through that surface is $\prod_{p}W_{p}=+1$ identically; this reduces the number of independent plaquettes by one. The flux through any other contractible loop is then the product of the fluxes through the plaquettes contained in the loop. Additionally, the surface on which we embed the graph may have noncontractible loops $\ell$ and there is an additional flux $W_{\ell}$ along each such loop (of which there are $b_{1}$—the first Betti number of the surface). There is also a weak symmetry $\mathcal{W}_{q}:=\mathcal{W}^{\mathrm{ch}}_{\langle ij\rangle}$ along each of the edges of the graph; in a doubled (i.e., vectorized) representation that we discuss in Sec. IV.2, the weak symmetry corresponds to the flux through an interlayer plaquette $q$ [see also Fig. 1(c)]. Any string $\mathcal{W}_{\mathcal{P}}^{\mathrm{ch}}$ is then given by an appropriate product of elementary fluxes $\mathcal{W}_{q}$. As detailed in App. B, the fluxes $\widetilde{W}_{p}$ and $\widetilde{W}_{\ell}$ are not independent as they are constrained by $W_{p}$, $W_{\ell}$ and $\mathcal{W}_{q}$. In total, we have $12E-N+$ 1 independent conserved fluxes: $E-N+1$ strong fluxes $W_{p}$ and $W_{\ell}$ and $E$ weak fluxes $\mathcal{W}_{q}$. The fluxes have eigenvalues ±$\pm 1$ (±$\pm\mathrm{i}$) if the length of their boundary is even (odd), as follows from the Hermiticity of the gamma matrices and the prefactor $\mathrm{i}^{|\mathcal{P}|}$ in Eq. (6)—this choice makes the antiunitary symmetries of the model manifest [8] (see also Sec. VI). $\mathcal{W}_{q}$ always has an even number of edges and thus always squares to $+1$. We denote the eigenvalues of $W_{p}$, $W_{\ell}$, $\widetilde{W}_{p}$, $\widetilde{W}_{\ell}$, and $\mathcal{W}_{q}$ by $w_{p}$, $w_{\ell}$, $\widetilde{w}_{p}$, $\widetilde{w}_{\ell}$, and $\mathfrak{w}_{q}$, respectively. We further denote a strong-flux configuration (i.e., an assignment $w_{p}=\pm 1$ or ±$\pm\mathrm{i}$ and $w_{\ell}=\pm 1$ or ±$\pm\mathrm{i}$ for all $p$ and $\ell$) by $\bm{w}$ and a weak-flux configuration (i.e., an assignment $\mathfrak{w}_{q}=\pm 1$ for all $q$) by $\bm{\mathfrak{w}}$; there are $2^{E-N+1}$ strong-flux configurations and $2^{E}$ weak-flux configurations.
@@ III.1 Steady states
As discussed above, there is a steady state for each eigenvalue of a strong symmetry, i.e., for each strong-flux configuration. Hence, we have $2^{E-N+1}$ distinct steady states, which we can build explicitly.
@@
is the projector into the symmetry sector labeled by the strong-flux configuration $\bm{w}$ and $\mathcal{Z}=2^{N-1}$ is the dimension of the symmetry sectors. [^2] This can be shown as follows. The total Hilbert space dimension is $\prod_{i}2^{z(i)/2}=2^{\sum_{i}z(i)/2}=2^{E}$, where in the last step we used the handshaking lemma. Assuming all the sectors have the same dimension, we divide the total dimension by the number $2^{E-N+1}$ of distinct steady states to find $\mathcal{Z}=2^{N-1}$. We justify our assumption once we introduce Majorana fermions in Sec. IV. Using Kitaev-model terminology [8], we say there is a vortex at plaquette $p$ if $w_{p}=-1$ and, therefore, a steady state is labeled by its vortex pattern. Because of the commutativity of all fluxes, we have $\mathcal{W}_{q}(\rho_{\bm{w}})=\rho_{\bm{w}}$, as required from a steady state under a weak symmetry.
@@
Figure 3: Logical operators of the Lindbladian gamma-matrix model. (a) Two noncontractible loops $\ell_{1}$ and $\ell_{2}$ for a random graph on the torus and two noncontractible loops ~$\tilde{\ell}_{1}$ and ~$\tilde{\ell}_{2}$ on the dual graph. $\ell_{a}$ and ~$\tilde{\ell}_{b}$ intersect and odd (even) number of times if $a=b$ ($a\neq b$). The same construction holds for higher-genus surfaces and any graph. (b) Construction of the logical operations in terms of gamma matrices.
@@ III.2 Symmetries and order parameters
We view $\rho_{\bm{w}}$ as logical states and to perform logical operations, we define (see also Fig. 3):
@@
where $a\in\{1,\dots,b_{1}\}$ and ~$\tilde{\ell}_{a}$ is a path on the dual graph such that $\ell_{a}$ and ~$\tilde{\ell}_{b}$ intersect and odd (even) number of times if $a=b$ ($a\neq b$). The index $i$ of $\Gamma_{i}^{\alpha}$ in $\overline{Z}_{a}$ can be chosen as either endpoint of each edge of the original graph crossed by ~$\tilde{\ell}_{a}$ and $\alpha=\alpha(\langle ij\rangle)$ is the color set by the Hamiltonian. These logical operators satisfy
$[\overline{X}_{a},\overline{X}_{b}]=[\overline{Z}_{a},\overline{Z}_{b}]=0\quad\text{and}\quad\overline{X}_{a}\overline{Z}_{b}=(-1)^{\delta_{ab}}\overline{Z}_{b}\overline{X}_{a}.$ (17)
The logical operators can be deformed without changing the commutation relations (17) by multiplying them with products of $K_{ij}$ and $\Gamma_{j}^{\mathrm{ch}}$. In particular, the two choices of index $i$ for $\Gamma_{i}^{\alpha}$ above differ by $K_{ij}$ on the edge $\langle ij\rangle$ crossed by ~$\tilde{\ell}_{a}$. The path of $\overline{X}_{a}$ is deformable by the multiplication by a $W_{p}$ (a direct-graph plaquette) adjacent to $\ell_{a}$; similarly, the path of $\overline{Z}_{a}$ on the dual graph is deformable by the multiplication by $\Gamma_{j}^{\mathrm{ch}}$ (a dual-graph plaquette) on a site $j$ adjacent to ~$\tilde{\ell}_{a}$. The logical operators act on $\rho_{\bm{w}}$ as
@@
with $\bm{w^{\prime}}$ such that $w^{\prime}_{p}=w_{p}$ for all $p$ and $w^{\prime}_{\ell_{b}}=-(-1)^{\delta_{ab}}w_{\ell_{b}}$.
@@ III.3 One-form strong-to-weak spontaneous symmetry breaking
If upon substituting $O\to O(x)\overline{O}(y)$, where $O(x)$ and $\overline{O}(y)$ are operators with charge $q_{O}$ and $-q_{O}$ and localized near $x$ and $y$, respectively, the correlators acquire a finite value when $|x-y|\to\infty$, then the respective symmetry is spontaneously broken by $\rho$ [55, 56, 57, 58, 59, 63, 64, 62]. More concretely, if the standard correlator (which can be viewed as a Rényi-1 correlator)
@@
for any $x\neq y$, where $\overline{Z}(x)$ is defined on a path that can be continuously deformed to ~$\tilde{\ell}_{a}$ and passes through site $x$. To arrive at Eq. (22) we used that $\overline{Z}_{a}(x)$ and $\overline{Z}_{a}(y)$ differ by a string of $K_{ij}$ and $\Gamma_{j}^{\mathrm{ch}}$ and the product of a steady state and any string $R$ of gamma matrices that is algebraically-independent of $W_{p}$ and $W_{\ell}$ (hence effectively acts as a tensor product $\rho_{\bm{w}}\otimes R$) is traceless. In Eq. (23), the string of gamma matrices arising from the left action commutes with $\rho_{\bm{w}}$ and cancels the string resulting from the right action; the numerator thus equals $\mathrm{Tr}[\rho_{\bm{w}}^{2}]$ and the result follows. Because the charged operators are extended one-dimensional objects, the noncontractible fluxes are one-form symmetries [133, 134, 135, 136] and the steady states exemplify one-form SW-SSB [59].
@@ III.4 Classical memory
While $\rho_{\bm{w}}$ is a logical state, it is only a classical memory, encoding $b_{1}$ classical bits. Indeed, an initial state with finite overlaps $\pi_{\bm{w}}$ with steady states $\rho_{\bm{w}}$ (where $\sum_{\bm{w}}\pi_{\bm{w}}=1$), evolves, at infinitely-long times, into the statistical mixture (i.e., incoherent superposition) $\sum_{\bm{w}}\pi_{\bm{w}}\rho_{\bm{w}}$. This idea can be formalized [50, 51] by identifying $\rho_{\bm{w}}$ as isolated extremal points in the convex set of steady states.
@@ IV.1 Majorana representation
The Lindbladian (10) becomes exactly solvable (i.e., reducible to exponentially many single-particle Lindbladians) if we fractionalize the degrees of freedom. To this end, we introduce $22k+$ 2 flavours of Majorana fermions per site, $b^{\mu}_{j}$ ($\mu\in\{1,\dots,2k\}$), $b_{j}^{\mathrm{ch}}=b_{j}^{2k+1}$, and $c_{j}$, and write the gamma matrices as
@@ IV.2 Vectorization
To proceed, we represent the Lindbladian superoperator as an operator on the space of density matrices, a procedure known as vectorization [76]. The vectorized density matrix is given by $|\rho\rangle=\sum_{mn}\langle m|\rho|n\rangle|m\rangle|n\rangle$, where $\{|n\rangle\}$ forms a basis of the original Hilbert space and $\{|m\rangle|n\rangle\}$ a basis of the doubled (vectorized) Hilbert space. The identity matrix is mapped to the (unnormalized) infinite-temperature thermofield double state $|0\rangle:=\sum_{m}|m\rangle|m\rangle$. For each Majorana fermion $c_{j}$ ($b^{\mu}_{j}$) in the original Hilbert space, we introduce two Majoranas in the doubled Hilbert space, $c_{j}$ and $\tilde{c}_{j}$ ($b^{\mu}_{j}$ and $\tilde{b}^{\mu}_{j}$); the former represent left action on the density matrix, the latter (a phase times) right action. When acting on the identity matrix, left and right multiplication should coincide, leading to $c_{j}|0\rangle=e^{\mathrm{i}\theta}\tilde{c}_{j}|0\rangle$. The phase $e^{\mathrm{i}\theta}$ is fixed by demanding anticommutation of Majoranas from different copies: acting on both sides of the equality with $c_{j}$ we have $|0\rangle=-e^{\mathrm{i}\theta}\tilde{c}_{j}c_{j}|0\rangle=-e^{2\mathrm{i}\theta}|0\rangle$, whence $e^{\mathrm{i}\theta}=\pm\mathrm{i}$. The choice of sign is arbitrary and we fix
@@
Using Eq. (27), one can then vectorize the left and right action of $b^{\mu}_{j}$ ($c_{j}$) on an arbitrary operator as follows [76]. For $M$ a fermion monomial with $DM=(-1)^{\eta_{M}}MD$, $D=\prod_{j}D_{j}$, we map $c_{j}M\to c_{j}|M\rangle$ and $Mc_{j}\to(-1)^{\eta_{M}}(-\mathrm{i}\tilde{c}_{j})|M\rangle$ ($b^{\mu}_{j}M\to b^{\mu}_{j}|M\rangle$ and $Mb^{\mu}_{j}\to(-1)^{\eta_{M}}(-\mathrm{i}\tilde{b}^{\mu}_{j})|M\rangle$). [^3] The inclusion of the fermionic parity $(-1)^{\eta_{M}}$ ensures that the product $c_{i}Mc_{j}$ maps to the same expression regardless of whether the left or right action is vectorized first: $(c_{i}M)c_{j}=(-1)^{\eta_{c_{i}M}}(-\mathrm{i}\tilde{c}_{j})|c_{j}M\rangle=(-1)^{\eta_{c_{i}M}}(-\mathrm{i}\tilde{c}_{j})c_{i}|M\rangle=(-1)^{\eta_{M}}c_{i}(-\mathrm{i}\tilde{c}_{j})|M\rangle=c_{i}(Mc_{j})$, where in the first step we used $|c_{j}M\rangle=c_{j}|M\rangle$ and in the second the anticommutation of $c_{i}$ and $\tilde{c}_{j}$ and $(-1)^{\eta_{c_{i}M}}=-(-1)^{\eta_{M}}$. A gamma matrix acting from the right gets mapped to $\widetilde{\Gamma}^{\mu}_{j}=-\mathrm{i}\tilde{b}^{\mu}_{j}\tilde{c}_{j}$ [note the relative minus sign compared with Eq. (24)]. With this mapping, the Lindbladian reads
@@
where the edge operators $u_{ij}$, $\tilde{u}_{ij}$, and $v_{j}$ are defined as

### v2

@@ II.5 Strong and weak symmetries
Unitary symmetries of open quantum systems come in two flavors: strong and weak [124] (sometimes also called exact and average, respectively), the properties of which we now briefly review. Let $S$ be a unitary operator in the physical Hilbert space. $S$ is a weak symmetry of $\mathcal{L}$ if $\mathcal{L}(S\rho S^{\dagger})=S\mathcal{L}(\rho)S^{\dagger}$, for any operator $\rho$. Introducing the superoperator $\mathcal{S}$ that acts by conjugation with $S$ (i.e., the adjoint action of $S$ on the space of operators), $\mathcal{S}(\rho)=S\rho S^{\dagger}$, a weak symmetry is equivalent to the superoperator commutation relation $\mathcal{L},\mathcal{S}]=0$. On the other hand, $S$ is a strong symmetry if $S$ commutes with all elements of the set $\{H,L_{j},L_{j}^{\dagger}\}$ ($j\in\{1,\dots,N\}$). It follows that the Lindbladian superoperator commutes with both $\mathcal{S}_{+}$ and $\mathcal{S}_{-}$, where $\mathcal{S}_{+}$ is the left action of $S$, $\mathcal{S}_{+}(\rho)=S\rho$, and $\mathcal{S}_{-}$ is the right action, $\mathcal{S}_{-}(\rho)=\rho S^{\dagger}$. Clearly, a strong symmetry is also weak, but the converse is not true. A simple example of a weak symmetry that is not strong (and that occurs in our model) is if $S$ commutes with $H$ but anticommutes with (some of) the jump operators.
@@ II.6 Symmetries of the gamma-matrix Lindbladian
Not all $W_{\mathcal{C}}$, $\widetilde{W}_{\mathcal{C}}$, and $\mathcal{W}_{\mathcal{P}}^{\mathrm{ch}}$ are independent. As worked out in detail in Appendix B, a minimal set—which labels the symmetry sectors—is $\{W_{p},W_{\ell},\mathcal{W}_{q}\}_{p,\ell,q}$. Here, $W_{p}$ are $W_{\mathcal{C}}$ for $\mathcal{C}=p$, with $p$ the elementary plaquettes (or faces) of the graph. If the graph is embedded on a closed surface, the total flux through that surface is $\prod_{p}W_{p}=+1$ identically; this reduces the number of independent plaquettes by one. The flux through any other contractible loop is then the product of the fluxes through the plaquettes contained in the loop. Additionally, the surface on which we embed the graph may have noncontractible loops $\ell$ and there is an additional flux $W_{\ell}$ along each such loop (of which there are $b_{1}$—the first Betti number of the surface). There is also a weak symmetry $\mathcal{W}_{q}:=\mathcal{W}^{\mathrm{ch}}_{\langle ij\rangle}$ along each of the edges of the graph; in a doubled (i.e., vectorized) representation that we discuss in Sec. IV.2, the weak symmetry corresponds to the flux through an interlayer plaquette $q$ [see also Fig. 1(c)]. Any string $\mathcal{W}_{\mathcal{P}}^{\mathrm{ch}}$ is then given by an appropriate product of elementary fluxes $\mathcal{W}_{q}$. As detailed in Appendix B, the fluxes $\widetilde{W}_{p}$ and $\widetilde{W}_{\ell}$ are not independent as they are constrained by $_{p}, WW_{\ell}$ and $\mathcal{W}_{q}$. In total, we have 2EN+12E-N+1 independent conserved fluxes: EN+1E-N+1 strong fluxes $W_{p}$ and $W_{\ell}$ and $E$ weak fluxes $\mathcal{W}_{q}$. The fluxes have eigenvalues $\pm 1$ ($\pm\mathrm{i}$) if the length of their boundary is even (odd), as follows from the Hermiticity of the gamma matrices and the prefactor $\mathrm{i}^{|\mathcal{P}|}$ in Eq. (6)—this choice makes the antiunitary symmetries of the model manifest [8] (see also Sec. VI). $\mathcal{W}_{q}$ always has an even number of edges and thus always squares to +1+1. We denote the eigenvalues of $_{p}, WW_{\ell}$, $\widetilde{W}_{p}$, $\widetilde{W}_{\ell}$, and $\mathcal{W}_{q}$ by $_{p}, ww_{\ell}$, $\widetilde{w}_{p}$, $\widetilde{w}_{\ell}$, and $\mathfrak{w}_{q}$, respectively. We further denote a strong-flux configuration (i.e., an assignment $w_{p}=\pm 1$ or $\pm\mathrm{i}$ and $w_{\ell}=\pm 1$ or $\pm\mathrm{i}$ for all $p$ and $\ell$) by $\bm{w}$ and a weak-flux configuration (i.e., an assignment $\mathfrak{w}_{q}=\pm 1$ for all $q$) by $\bm{\mathfrak{w}}$; there are $2^{E-N+1}$ strong-flux configurations and $2^{E}$ weak-flux configurations.
@@ III Mixed-state topological order
Figure 3: Logical operators of the Lindbladian gamma-matrix model. (a) Two noncontractible loops $\ell_{1}$ and $\ell_{2}$ for a random graph on the torus and two noncontractible loops $\tilde{\ell}_{1}$ and $\tilde{\ell}_{2}$ on the dual graph. $\ell_{a}$ and $\tilde{\ell}_{b}$ intersect and odd (even) number of times if $a=b$ ($a\neq b$). The same construction holds for higher-genus surfaces and any graph. (b) Construction of the logical operations in terms of gamma matrices.
@@ III.1 Steady states
The steady states $\rho_{\mathrm{ss}}$ of $\mathcal{L}$ satisfy $\dot{\rho}_{\mathrm{ss}}=\mathcal{L}(\rho_{\mathrm{ss}})=0$, i.e., they are eigenoperators of $\mathcal{L}$ with zero eigenvalue. As discussed above, there is one such steady state for each eigenvalue of a strong symmetry, i.e., for each strong-flux configuration. Hence, we have $2^{E-N+1}$ distinct steady states, which we can build explicitly.
@@
is the projector into the symmetry sector labeled by the strong-flux configuration $\bm{w}$, and $\mathcal{Z}=2^{N-1}$ is the dimension of the symmetry sectors. [^2] This can be shown as follows. The total Hilbert space dimension is $\prod_{i}2^{z(i)/2}=2^{\sum_{i}z(i)/2}=2^{E}$, where in the last step we used the handshaking lemma. Assuming all the sectors have the same dimension, we divide the total dimension by the number $2^{E-N+1}$ of distinct steady states to find $\mathcal{Z}=2^{N-1}$. We justify our assumption once we introduce Majorana fermions in Sec. IV. Using Kitaev-model terminology [8], we say there is a vortex at plaquette $p$ if $w_{p}=-1$ and, therefore, a steady state is labeled by its vortex pattern. Because of the commutativity of all fluxes, we have $\mathcal{W}_{q}(\rho_{\bm{w}})=\rho_{\bm{w}}$, as required from a steady state under a weak symmetry.
@@ III.2 Symmetries and order parameters
We view $\rho_{\bm{w}}$ as logical states, and to perform logical operations, we define (see also Fig. 3):
@@
where $a\in\{1,\dots,b_{1}\}$ and $\tilde{\ell}_{a}$ is a path on the dual graph such that $\ell_{a}$ and $\tilde{\ell}_{b}$ intersect an odd (even) number of times if $a=b$ ($a\neq b$). The index $i$ of $\Gamma_{i}^{\alpha}$ in $\overline{Z}_{a}$ can be chosen as either endpoint of each edge of the original graph crossed by $\tilde{\ell}_{a}$, and $\alpha=\alpha(\langle ij\rangle)$ is the color set by the Hamiltonian. These logical operators satisfy
$\overline{X}_{a},\overline{X}_{b}]=[\overline{Z}_{a},\overline{Z}_{b}]=0\quad\text{and}\quad\overline{X}_{a}\overline{Z}_{b}=(-1)^{\delta_{ab}}\overline{Z}_{b}\overline{X}_{a}.$ (17)
The logical operators can be deformed without changing the commutation relations (17) by multiplying them with products of $K_{ij}$ and $\Gamma_{j}^{\mathrm{ch}}$. In particular, the two choices of index $i$ for $\Gamma_{i}^{\alpha}$ above differ by $K_{ij}$ on the edge $\langle ij\rangle$ crossed by $\tilde{\ell}_{a}$. The path of $\overline{X}_{a}$ is deformable by the multiplication by a $W_{p}$ (a direct-graph plaquette) adjacent to $\ell_{a}$; similarly, the path of $\overline{Z}_{a}$ on the dual graph is deformable by the multiplication by $\Gamma_{j}^{\mathrm{ch}}$ (a dual-graph plaquette) on a site $j$ adjacent to $\tilde{\ell}_{a}$. The logical operators act on $\rho_{\bm{w}}$ as
@@
with $\bm{w^{\prime}}$ such that $w^{\prime}_{p}=w_{p}$ for all $p$ and $w^{\prime}_{\ell_{b}}=(-1)^{\delta_{ab}}w_{\ell_{b}}$.
@@ III.3 One-form strong-to-weak spontaneous symmetry breaking
If upon substituting $O\to O(x)\overline{O}(y)$, where $O(x)$ and $\overline{O}(y)$ are operators with charge $q_{O}$ and $q_{O}$ and localized near $x$ and $y$, respectively, the correlators acquire a finite value when $|x-y|\to\infty$, then the respective symmetry is spontaneously broken by $\rho$ [56, 57, 58, 59, 60, 64, 65, 63]. More concretely, if the standard correlator (which can be viewed as a Rényi-1 correlator)
@@
for any $x\neq y$, where $\overline{Z}(x)$ is defined on a path that can be continuously deformed to $\tilde{\ell}_{a}$ and passes through site $x$. To arrive at Eq. (22) we used that $\overline{Z}_{a}(x)$ and $\overline{Z}_{a}(y)$ differ by a string of $K_{ij}$ and $\Gamma_{j}^{\mathrm{ch}}$, and the product of a steady state and any string $R$ of gamma matrices that is algebraically independent of $W_{p}$ and $W_{\ell}$ (hence effectively acts as a tensor product $\rho_{\bm{w}}\otimes R$) is traceless. In Eq. (23), the string of gamma matrices arising from the left action commutes with $\rho_{\bm{w}}$ and cancels the string resulting from the right action; the numerator thus equals $\mathrm{Tr}[\rho_{\bm{w}}^{2}]$ and the result follows. Because the charged operators are extended one-dimensional objects, the noncontractible fluxes are one-form symmetries [136, 137, 138, 139] and the steady states exemplify one-form SW-SSB [60].
@@ III.4 Classical memory
While $\rho_{\bm{w}}$ is a logical state, it is only a classical memory, encoding $b_{1}$ classical bits. Indeed, an initial state with finite overlaps $\pi_{\bm{w}}$ with the steady states $\rho_{\bm{w}}$ (where $\sum_{\bm{w}}\pi_{\bm{w}}=1$), evolves, at infinitely-long times, into the statistical mixture (i.e., incoherent superposition) $\sum_{\bm{w}}\pi_{\bm{w}}\rho_{\bm{w}}$. This idea can be formalized [51, 52] by identifying $\rho_{\bm{w}}$ as isolated extremal points in the convex set of steady states.
@@ IV.1 Majorana representation
The Lindbladian (10) becomes exactly solvable (i.e., reducible to exponentially many single-particle Lindbladians) if we fractionalize the degrees of freedom. To this end, we introduce $2k+2$ flavours of Majorana fermions per site, $b^{\mu}_{j}$ ($\mu\in\{1,\dots,2k\}$), $b_{j}^{\mathrm{ch}}=b_{j}^{2k+1}$, and $c_{j}$, and write the gamma matrices as
@@ IV.2 Vectorization
To proceed, we represent the Lindbladian superoperator as an operator on the space of density matrices, a procedure known as vectorization [77]. The vectorized density matrix is given by $|\rho\rangle=\sum_{mn}\langle m|\rho|n\rangle|m\rangle|n\rangle$, where $\{|n\rangle\}$ forms a basis of the original Hilbert space and $\{|m\rangle|n\rangle\}$ is a basis of the doubled (vectorized) Hilbert space. The identity matrix is mapped to the (unnormalized) infinite-temperature thermofield double state $|0\rangle:=\sum_{m}|m\rangle|m\rangle$. For each Majorana fermion $c_{j}$ ($b^{\mu}_{j}$) in the original Hilbert space, we introduce two Majoranas in the doubled Hilbert space, $c_{j}$ and $\tilde{c}_{j}$ ($b^{\mu}_{j}$ and $\tilde{b}^{\mu}_{j}$); the former represent left action on the density matrix, while the latter represent (a phase times) right action. When acting on the identity matrix, left and right multiplication should coincide, leading to $c_{j}|0\rangle=e^{\mathrm{i}\theta}\tilde{c}_{j}|0\rangle$. The phase $e^{\mathrm{i}\theta}$ is fixed by demanding anticommutation of Majoranas from different copies: acting on both sides of the equality with $c_{j}$ we have $|0\rangle=-e^{\mathrm{i}\theta}\tilde{c}_{j}c_{j}|0\rangle=-e^{2\mathrm{i}\theta}|0\rangle$, whence $e^{\mathrm{i}\theta}=\pm\mathrm{i}$. The choice of sign is arbitrary and we fix
@@
Using Eq. (27), one can then vectorize the left and right action of $b^{\mu}_{j}$ ($c_{j}$) on an arbitrary operator as follows [77]. For $M$ a fermion monomial with $DM=(-1)^{\eta_{M}}MD$, $D=\prod_{j}D_{j}$, we map $c_{j}M\to c_{j}|M\rangle$ and $Mc_{j}\to(-1)^{\eta_{M}}(-\mathrm{i}\tilde{c}_{j})|M\rangle$ ($b^{\mu}_{j}M\to b^{\mu}_{j}|M\rangle$ and $Mb^{\mu}_{j}\to(-1)^{\eta_{M}}(-\mathrm{i}\tilde{b}^{\mu}_{j})|M\rangle$). [^3] The inclusion of the fermionic parity $(-1)^{\eta_{M}}$ ensures that the product $c_{i}Mc_{j}$ maps to the same expression regardless of whether the left or right action is vectorized first: $(c_{i}M)c_{j}=(-1)^{\eta_{c_{i}M}}(-\mathrm{i}\tilde{c}_{j})|c_{i}M\rangle=(-1)^{\eta_{c_{i}M}}(-\mathrm{i}\tilde{c}_{j})c_{i}|M\rangle=(-1)^{\eta_{M}}c_{i}(-\mathrm{i}\tilde{c}_{j})|M\rangle=c_{i}(Mc_{j})$, where in the second equality we used $|c_{j}M\rangle=c_{j}|M\rangle$ and in the third the anticommutation of $c_{i}$ and $\tilde{c}_{j}$ and $(-1)^{\eta_{c_{i}M}}=-(-1)^{\eta_{M}}$. A gamma matrix acting from the right gets mapped to $\widetilde{\Gamma}^{\mu}_{j}=-\mathrm{i}\tilde{b}^{\mu}_{j}\tilde{c}_{j}$ [note the relative minus sign compared with Eq. (24)]. With this mapping, the Lindbladian reads
@@
where the edge operators (referred to as gauge fields) $u_{ij}$, $\tilde{u}_{ij}$, and $v_{j}$ are defined as